OpenStudy (anonymous):

you purchase a five-year savings bond. The Value of the bond increases with Simple Interest by 5% each year. After five years, what is the total percentage increase from the bond's initial value? (a) 25.0% (b) 25.8% (c) 27.6% (d) 28.4% (e) None of the above

OpenStudy (anonymous):

first yr: 100% x 105/100 = 105% second year: 105% x 105/100 = 110.25% third yr: 110.25% x 105/100 = ... etc... Do you understand it? :)

OpenStudy (anonymous):

yup, thank you very much

OpenStudy (anonymous):

in simple interest the interest earned is not taken into account